# MnLargeSymbols’164 MnLargeSymbols’171

### v2

@@ 3.1.2 Properties of the value function
Step i) [mild form of HJB equation] Consider, for given $\phi\in C^{0}(\mathbb{R}\times\mathbb{R}_{+}$), the HJB equation (i.e. the first equation of (10) with generic final datum $\phi$):
$\begin{cases}-\partial_{t}V+\rho V=H_{1}(t,x,h,\mu(t),D_{h}V)+\frac{1}{2}\chi^{2}h^{2}D^{2}_{hh}V+\frac{1}{2}\epsilon^{2}D^{2}_{xx}V+H_{0}(D_{x}V)&\mbox{ in}(0,T)\times\mathbb{R}\times\mathbb{R}_{++}\\ V(T,x,h)=\phi(x,h)\quad\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}\\ V(t,x,0)=0\quad\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (25)
By reversing time, i.e. calling $v(t,x,h):=V(T-t,x,h)$ this PDE becomes
$\begin{cases}\partial_{t}v=\frac{1}{2}\epsilon^{2}D^{2}_{xx}v+\frac{1}{2}\chi^{2}h^{2}D^{2}_{hh}v+\rho v+H_{0}(D_{x}v)+H_{1}(T-t,x,h,\mu(T-t),D_{h}v)&\mbox{in }(0,T)\times\mathbb{R}\times\mathbb{R}_{++}\\ v(0,x,h)=\phi(x,h)\quad\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}\\ v(t,x,0)=0\quad\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (26)
@@
$\begin{cases}\partial_{t}w=\frac{1}{2}\epsilon^{2}D^{2}_{xx}w+\frac{1}{2}\chi^{2}D^{2}_{yy}w-\frac{1}{2}\chi D_{y}w+\rho w+H_{0}(D_{x}w)+H_{1}(T-t,x,e^{y},\mu(T-t),e^{-y}D_{y}w)&\mbox{ in }(0,T)\times\mathbb{R}\times\mathbb{R}\\ w(0,x,y)=\phi(x,e^{y})\quad\mbox{ in }\mathbb{R}\times\mathbb{R}\\ w(t,x,-\infty)=0\quad\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (27)
@@
where the weight is chosen in way that the value function, after the change of variable, belongs to that space. We can then apply the results of [27, Chapter 7], more precisely Theorem 7.1.3 and Proposition 7.2.1. Such results establish that, under suitable assumptions on the nonlinearities the above PDE (27) has a unique mild solution satisfying the estimate
@@
It then remains to check that the nonlinearity here satisfies the assumptions of Theorem 7.1.3 and Proposition 7.2.1 of [27, Chapter 7]. This is immediate consequence of the Remark 2.4 (for the part of $H_{0}$) and of the subsequent Lemma 3.4 (for the part of $H_{1}$). ∎
@@
=$=\begin{cases}u_{\sigma}\left(A(x)f(h)^{1-\gamma}F(x,\mu)^{\gamma}\right)-\zeta hp&\text{ for }p\leq p_{0}\\ \left(1-\left(\frac{p_{0}}{p}\right)^{\frac{1}{1-\eta}}\right)f(h)F(x,\mu)p-\zeta hp+u_{\sigma}\left(A(x)f(h)^{1-\gamma}F(x,\mu)^{\gamma}\right)\left(\frac{p_{0}}{p}\right)^{\frac{\eta}{1-\eta}}&\text{ for }p>p_{0}\end{cases}$ (28)
=$=\begin{cases}\frac{1}{1-\sigma}A(x)^{1-\sigma}f(h)^{\eta}F(x,\mu)^{\gamma(1-\sigma)}-\zeta hp&\text{ for }p\leq p_{0}\\ \left(f(h)F(x,\mu)-\zeta h\right)p+\frac{1-\eta}{\eta}p_{0}^{\frac{1}{1-\eta}}f(h)F(x,\mu)p^{-\frac{\eta}{1-\eta}}&\text{ for }p>p_{0}.\end{cases}$ (29)
@@
$-\zeta h\leq D_{p}H_{1}(x,h,\mu,p)\leq\left(\zeta+2F(x,\mu)\right)h,$ (31)
@@
The function $D_{p}H_{1}$ is locally Lispchitz continuous for $M(\mu)$ away from 00, in the following sense: for every $N\geq 1$ there exists a positive constant $C_{N}$ such that for every $\mu_{1},\mu_{2}$ satisfying $\frac{1}{N}<M(\mu_{1})\wedge M(\mu_{2})\leq\left(M_{2}(\mu_{1})\vee M_{2}(\mu_{2})\right)^{\frac{1}{2}}<N$, for every $h_{1}\leq h_{2}<N$, for every $x_{1},x_{2}$ and for every $p_{1},p_{2}$ we have
$\left|D_{p}H_{1}(x_{1},h_{2},\mu_{1},p_{1})-D_{p}H_{1}(x_{2},h_{2},\mu_{2},p_{2})\right|\leq C_{N}\left(|x_{1}-x_{2}|+|h_{1}-h_{2}|+\mathcal{W}_{1}(\mu_{1},\mu_{2})+|p_{1}-p_{2}|\right).$
@@
otherwise. Then $p>p_{0}(x,h,\mu)$ if and only if $h>h_{0}(x,\mu,p)$, and we have $p=p_{0}(x,\mu,h_{0}(x,\mu,p))$ whenever $p\neq 0$ and $M(\mu)>0$. By (30) and the definitions of $p_{0}$ and $h_{0}$ we have
@@
otherwise, so that $p>p_{0}(x,h,\mu)$ if and only if $F(x,\mu)>F_{0}(x,h,p)$, and we have $F_{0}\left(x,h,p_{0}(x,h,\mu)\right)=F(x,\mu)$ whenever $p$ and $h$ are nonzero. Consider $\mu_{1},\mu_{2}$ such that $0<\underline{M}:=M_{2}(\mu_{1})\wedge M_{2}(\mu_{2})$, and set also $\overline{M}_{2}:=M_{2}(\mu_{1})\vee M_{2}(\mu_{2})$, $\alpha=\frac{1-\gamma(1-\sigma)}{1-\eta}$ and $\beta=\frac{1-\sigma}{1-\eta}$. If $F_{0}(x,h,p)>F(x,\mu_{1})\vee F(x,\mu_{2})$ then
@@
If $F_{0}(x,h,p)\leq F(x,\mu_{1})\wedge F(x,\mu_{2})$ we get, using (30), Lemma 3.1, the fact that $p\mapsto p^{-\frac{1}{1-\eta}}$ is decreasing and $11-\alpha<$ 1,
@@
$\phantom{aaaaaaaaaa}+p_{0}(x,h,\mu_{1})^{-\frac{1}{1-\eta}}(1-\gamma)^{\frac{1}{1-\eta}}A(x)^{\beta}\left|1-\alpha\right|\left(\frac{\theta}{\Theta}\underline{M}\right)^{-\alpha}\left|F(x,\mu_{1})-F(x,\mu_{2})\right|$
@@
$\leq\frac{\Theta}{\theta}M(\mu_{1})p_{0}(x,h,\mu_{1})^{-\frac{1}{1-\eta}}(1-\gamma)^{\frac{1}{1-\eta}}A(x)^{\beta}\left|\gamma(1-\sigma)-1\right|\left(\frac{\theta}{\Theta}\underline{M}\right)^{\gamma(1-\sigma)-2}\left|F(x,\mu_{1})-F(x,\mu_{1})\right|$
$\leq\frac{\Theta}{\theta}M(\mu_{1})f(h)F(x,\mu_{1})^{\alpha}\left(1-\gamma(1-\sigma)\right)\left(\frac{\theta}{\Theta}\underline{M}\right)^{\gamma(1-\sigma)-2}\frac{\Theta}{\theta^{2}}\left(M_{2}(\mu_{1})^{\frac{1}{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}\right)+\Theta\right)\mathcal{W}_{2}(\mu_{1},\mu_{2})$
$\leq\left(\frac{\Theta}{\theta}\overline{M}_{2}^{\frac{1}{2}}\right)^{1+\alpha}L_{f}h(1-\gamma(1-\sigma))\left(\frac{\Theta}{\theta}\frac{1}{\underline{M}}\right)^{2-\gamma(1-\sigma)}\frac{\Theta}{\theta^{2}}\left(M_{2}(\mu_{1})^{\frac{1}{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}\right)+\Theta\right)\mathcal{W}_{2}(\mu_{1},\mu_{2}).$
@@
so that $G_{0}\left(h,p_{0}(x,h,\mu)\right)=A(x)^{1-\sigma}F(x,\mu)^{\gamma(1-\sigma)-1}$ and $A(x_{1})^{1-\sigma}F(x_{1},\mu)^{\gamma(1-\sigma)-1}<G_{0}(h,p)\leq A(x_{2})^{1-\sigma}F(x_{2},\mu)^{\gamma(1-\sigma)-1}$. Then
[…]
$\leq f(h)\frac{\Theta}{\theta}M(\mu)p_{0}(x_{1},h,\mu)^{-\frac{1}{1-\eta}}(1-\gamma)^{\frac{1}{1-\eta}}f(h)^{-1}\left(G_{0}(h,p)^{\frac{1}{1-\eta}}-\left(A(x_{1})^{1-\sigma}F(x_{1},\mu)^{\gamma(1-\sigma)-1}\right)^{\frac{1}{1-\eta}}\right)$
$=f(h)F(x_{1},\mu)\left(A(x_{1})^{1-\sigma}F(x_{1},\mu)^{\gamma(1-\sigma)-1}\right)^{-\frac{1}{1-\eta}}\left(G_{0}(h,p)^{\frac{1}{1-\eta}}-G_{0}\left(h,p_{0}(x_{1},h,\mu)\right)^{\frac{1}{1-\eta}}\right)$
@@
$\leq L_{f}h\left(\frac{\Theta}{\theta}M(\mu)\right)^{1+\alpha}\underline{A}^{-\beta}\left(F(x_{1},\mu)^{-\alpha}\left|A(x_{1})^{\beta}-A(x_{2})^{\beta}\right|+A(x_{2})^{\beta}\left|F(x_{1},\mu)^{-\alpha}-F(x_{2},\mu)^{-\alpha}\right|\right)$
@@
$\begin{cases}\partial_{t}\mu(t)=\frac{1}{2}\epsilon^{2}D^{2}_{xx}\mu(t)+\frac{1}{2}\chi^{2}D^{2}_{hh}\left(h^{2}\mu(t)\right)-D_{x}\left(D_{p}H_{0}(D_{x}V(t,x,h))\mu(t)\right)-D_{h}\left(D_{p}H_{1}(x,h,\mu(t),D_{h}V(t,x,h))\mu(t)\right),\\ \mu(0)=\mu_{0}\end{cases}$ (34)
@@
There exists $T>0$ such that we have existence of a strong solution $(\tilde{x}(\cdot),\tilde{h}(\cdot))$ to (35) on $[0,T]$, which is morevoer pathwise unique.
@@
$\left|\begin{pmatrix}D_{p}H_{0}\left(D_{x}V(t,x_{1},h_{1})\right)-D_{p}H_{0}\left(D_{x}V(t,x_{2},h_{2})\right)\\ D_{p}H_{1}\left(x_{1},h_{1},\mu(t),D_{h}V(t,x_{1},h_{1})\right)-D_{p}H_{1}\left(x_{2},h_{2},\mu(t),D_{h}V(t,x_{2},h_{2})\right)\end{pmatrix}\right|\leq L_{H,N}\left|(x_{1}-x_{2},h_{1}-h_{2})\right|.$
@@
[…]
@@
[…]
@@
[…]
If $10<\mu_{0}\left(\mathbb{R}\times\{0\}\right)<$ 1 (i.e. if $10<\mathbb{P}(h_{0}=0)<$ 1), we can consider the two conditional laws of $(\tilde{x}(\cdot),\tilde{h}(\cdot))$ with respect to the sets $\{h_{0}=0\}$ and $\{h_{0}>0\}$. By the arguments above, the measure defined by
@@
for every measurable $A\subset C([0,+\infty))$ gives a weak solution to (36); moreover the process $(\hat{x}(\cdot),\hat{h}(\cdot))$ with law $\mathscr{L}$ is pathwise unique, therefore the solution is strong by the Yamada-Watanabe theorem.
This shows that, for every $\mu(\cdot)\in C^{\frac{1}{2}}\left([0,T_{1}];\mathcal{P}_{2}\right)$ such that $\mu(0)=\mu_{0}$ and every $\mathbb{R}\times\mathbb{R}_{+}$-valued random variable $(x_{0},h_{0})$ with law $\mu_{0}$, equation (36) has a unique strong solution on $[0,T_{1}]$.
@@
$\mathcal{Q}_{K_{1},K_{2}}=\left\{\mu(\cdot)\colon[0,T]\rightarrow\mathcal{P}_{2}\text{ s.t. }\mu(0)=\mu_{0},\sup_{t\in[0,T]}P_{2}(\mu(t))\leq 2K_{1},\sup_{\begin{subarray}{c}s,t\in[0,T]\\ s\neq t\end{subarray}}\frac{\mathcal{W}_{2}(\mu(s),\mu(t))}{|s-t|^{\frac{1}{2}}}\leq K_{2}\right\}.$ (39)
@@
Considering only the equation for $\hat{h}(\cdot)$, we have
$\mathbb{E}\left[\sup_{t\in[0,T]}\hat{h}(t)^{2}\right]$ $\leq 3\mathbb{E}[h_{0}^{2}]+3Tg_{1}^{2}(\overline{M}(\mu(\cdot)))\int_{0}^{T}\mathbb{E}\left[\sup_{s\in[0,t]}h(s)^{2}\right]\mathrm{d}t+3\chi^{2}\mathbb{E}\left[\sup_{t\in[0,T]}\left|\int_{0}^{t}h(s)\mathrm{d}W(s)\right|^{2}\right]$
@@
$\leq 2\mathbb{E}\left[\left|\int_{0}^{t}\left(D_{p}H_{1}\left(\hat{x}^{\mu}(s),\hat{h}^{\mu}(s),\mu(s),D_{h}V\left(s,\hat{x}^{\mu}(s),\hat{h}^{\mu}(s)\right)\right)-D_{p}H_{1}\left(\hat{x}^{\nu}(s),\hat{h}^{\nu}(s),\mu(s),D_{h}V\left(s,\hat{x}^{\nu}(s),\hat{h}^{\nu}(s)\right)\right)\right)\mathrm{d}s\right|^{2}\right]$
@@
If instead $10<\mu_{0}(\mathbb{R}\times\{0\})<$ 1 then $M(\mu_{0})>0$ and we can again condition on the sets $\{h_{0}=0\}$ and $\{h_{0}>0\}$. As the solution $(\tilde{x}(\cdot),\tilde{h}(\cdot))$ is uniquely determined on each set, well-posedness of (35) follows. ∎
@@
×$\int_{\mathbb{R}\times\mathbb{R}_{+}}u(t,x,h)\mu(t,\mathrm{d}x,\mathrm{d}h)-\int_{\mathbb{R}\times\mathbb{R}_{+}}u(t_{0},x,h)\mu(s,\mathrm{d}x,\mathrm{d}h)=\int_{t_{0}}^{t}\int_{\mathbb{R}\times\mathbb{R}_{+}}\frac{\partial u}{\partial t}(r,x,h)\mu(r,\mathrm{d}x,\mathrm{d}h)\mathrm{d}r\\ +\frac{1}{2}\int_{t_{0}}^{t}\int_{\mathbb{R}\times\mathbb{R}_{+}}\frac{1}{2}\mathrm{Tr}\left[g(x,h)g^{\ast}(x,h)D^{2}u(r,x,h)\right]\mu(r,\mathrm{d}x,\mathrm{d}h)\mathrm{d}r\\ +\int_{t_{0}}^{t}\int_{\mathbb{R}\times\mathbb{R}_{+}}e^{\mu}(r,x,h)\cdot Du(r,x,h)\mu(r,\mathrm{d}x,\mathrm{d}h)$
@@
×$\int_{\mathbb{R}\times\mathbb{R}_{+}}\phi(x,h)\mu(t,\mathrm{d}x,\mathrm{d}h)=\int_{\mathbb{R}\times\mathbb{R}_{+}}u(0,x,h)\mu_{0}(\mathrm{d}x,\mathrm{d}h);$

### v3

@@ 3.1.2 Properties of the value function
Consider, for given $\phi\in C^{0}(\mathbb{R}\times\mathbb{R}_{+}$), the HJB equation (i.e. the first equation of (10) with generic final datum $\phi$):
$\begin{cases}-\partial_{t}V+\rho V=H_{1}(t,x,h,\mu(t),D_{h}V)+H_{0}(D_{x}V)&\\ \phantom{aaaaaaaaaaaa}+\frac{1}{2}\chi^{2}h^{2}D^{2}_{hh}V+\frac{1}{2}\epsilon^{2}D^{2}_{xx}V&\mbox{ in }(0,T)\times\mathbb{R}\times\mathbb{R}_{++}\\ V(T,x,h)=\phi(x,h)&\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}\\ V(t,x,0)=0&\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (25)
By reversing time, i.e. calling v(t,x,h):=V(Tt,x,h)v(t,x,h):=V(T-t,x,h) this PDE becomes
$\begin{cases}\partial_{t}v=\rho v+H_{0}(D_{x}v)+H_{1}(T-t,x,h,\mu(T-t),D_{h}v)&\\ \phantom{aaaaaaaaa}+\frac{1}{2}\epsilon^{2}D^{2}_{xx}v+\frac{1}{2}\chi^{2}h^{2}D^{2}_{hh}v&\mbox{ in }(0,T)\times\mathbb{R}\times\mathbb{R}_{++}\\ v(0,x,h)=\phi(x,h)&\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}\\ v(t,x,0)=0&\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (26)
@@
$\begin{cases}\partial_{t}w=\rho w+H_{0}(D_{x}w)+H_{1}(T-t,x,e^{y},\mu(T-t),e^{-y}D_{y}w)&\\ \phantom{aaaaaaaaaa}+\frac{1}{2}\epsilon^{2}D^{2}_{xx}w+\frac{1}{2}\chi^{2}D^{2}_{yy}w-\frac{1}{2}\chi D_{y}w&\mbox{ in }(0,T)\times\mathbb{R}\times\mathbb{R}\\ w(0,x,y)=\phi(x,e^{y})&\mbox{ in }\mathbb{R}\times\mathbb{R}\\ w(t,x,-\infty)=0&\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (27)
@@
where the weight is chosen in way that the value function, after the change of variable, belongs to that space. We can then apply the results of [23, Chapter 7], more precisely Theorem 7.1.3 and Proposition 7.2.1. Such results establish that, under suitable assumptions on the nonlinearities the above PDE (27) has a unique mild solution satisfying the estimate
@@
It then remains to check that the nonlinearity here satisfies the assumptions of Theorem 7.1.3 and Proposition 7.2.1 of [23, Chapter 7]. This is immediate consequence of the Remark 2.4 (for the part of $H_{0}$) and of the subsequent Lemma 3.4 (for the part of $H_{1}$). ∎
@@
$=\begin{cases}u_{\sigma}\left(A(x)f(h)^{1-\gamma}F(x,\mu)^{\gamma}\right)-\zeta hp&\text{ for }p\leq p_{0}\\ \left(1-\left(\frac{p_{0}}{p}\right)^{\frac{1}{1-\eta}}\right)f(h)F(x,\mu)p-\zeta hp+u_{\sigma}\left(A(x)f(h)^{1-\gamma}F(x,\mu)^{\gamma}\right)\left(\frac{p_{0}}{p}\right)^{\frac{\eta}{1-\eta}}&\text{ for }p>p_{0}\end{cases}$ (28)
$=\begin{cases}\frac{1}{1-\sigma}A(x)^{1-\sigma}f(h)^{\eta}F(x,\mu)^{\gamma(1-\sigma)}-\zeta hp&\text{ for }p\leq p_{0}\\ \left(f(h)F(x,\mu)-\zeta h\right)p+\frac{1-\eta}{\eta}p_{0}^{\frac{1}{1-\eta}}f(h)F(x,\mu)p^{-\frac{\eta}{1-\eta}}&\text{ for }p>p_{0}.\end{cases}$ (29)
@@
$\zeta h\leq D_{p}H_{1}(x,h,\mu,p)\leq\left(\zeta+2F(x,\mu)\right)h,$ (31)
@@
The function $D_{p}H_{1}$ is locally Lispchitz continuous for $M(\mu)$ away from 0, in the following sense: for every $N\geq 1$ there exists a positive constant $C_{N}$ such that for every $\mu_{1},\mu_{2}$ satisfying $\frac{1}{N}<M(\mu_{1})\wedge M(\mu_{2})\leq\left(M_{2}(\mu_{1})\vee M_{2}(\mu_{2})\right)^{\frac{1}{2}}<N$, for every $h_{1}\leq h_{2}<N$, for every $x_{1},x_{2}$ and for every $p_{1},p_{2}$ we have
$\left|D_{p}H_{1}(x_{1},h_{2},\mu_{1},p_{1})-D_{p}H_{1}(x_{2},h_{2},\mu_{2},p_{2})\right|\\ \leq C_{N}\left(|x_{1}-x_{2}|+|h_{1}-h_{2}|+\mathcal{W}_{1}(\mu_{1},\mu_{2})+|p_{1}-p_{2}|\right).$
@@
otherwise. It follows that $p>p_{0}(x,h,\mu)$ if and only if $h>h_{0}(x,\mu,p)$, and we have $p=p_{0}(x,\mu,h_{0}(x,\mu,p))$ whenever $p\neq 0$ and $M(\mu)>0$. By (30) and the definitions of $p_{0}$ and $h_{0}$ we have
@@
otherwise. This implies that $p>p_{0}(x,h,\mu)$ if and only if $F(x,\mu)>F_{0}(x,h,p)$, and we have $F_{0}\left(x,h,p_{0}(x,h,\mu)\right)=F(x,\mu)$ whenever $p$ and $h$ are nonzero. Consider $\mu_{1},\mu_{2}$ such that $0<\underline{M}:=M_{2}(\mu_{1})\wedge M_{2}(\mu_{2})$, and set also $\overline{M}_{2}:=M_{2}(\mu_{1})\vee M_{2}(\mu_{2})$, $\alpha=\frac{1-\gamma(1-\sigma)}{1-\eta}$ and $\beta=\frac{1-\sigma}{1-\eta}$. If $F_{0}(x,h,p)>F(x,\mu_{1})\vee F(x,\mu_{2})$ then
@@
If $F_{0}(x,h,p)\leq F(x,\mu_{1})\wedge F(x,\mu_{2})$ we get, using (30), Lemma 3.1, the fact that $p\mapsto p^{-\frac{1}{1-\eta}}$ is decreasing and $1-\alpha<1$,
@@
$\phantom{aaaaa}+p_{0}(x,h,\mu_{1})^{-\frac{1}{1-\eta}}(1-\gamma)^{\frac{1}{1-\eta}}A(x)^{\beta}\left|1-\alpha\right|\left(\frac{\theta}{\Theta}\underline{M}\right)^{-\alpha}\left|F(x,\mu_{1})-F(x,\mu_{2})\right|$
@@
$\leq\frac{\Theta}{\theta}M(\mu_{1})p_{0}(x,h,\mu_{1})^{-\frac{1}{1-\eta}}(1-\gamma)^{\frac{1}{1-\eta}}A(x)^{\beta}$
$\phantom{aaaaaa}\cdot\left|\gamma(1-\sigma)-1\right|\left(\frac{\theta}{\Theta}\underline{M}\right)^{\gamma(1-\sigma)-2}\left|F(x,\mu_{1})-F(x,\mu_{1})\right|$
$\leq\frac{\Theta}{\theta}M(\mu_{1})f(h)F(x,\mu_{1})^{\alpha}\left(1-\gamma(1-\sigma)\right)$
$\phantom{aaaaaa}\cdot\left(\frac{\theta}{\Theta}\underline{M}\right)^{\gamma(1-\sigma)-2}\frac{\Theta}{\theta^{2}}\left(M_{2}(\mu_{1})^{\frac{1}{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}\right)+\Theta\right)\mathcal{W}_{2}(\mu_{1},\mu_{2})$
$\leq\left(\frac{\Theta}{\theta}\overline{M}_{2}^{\frac{1}{2}}\right)^{1+\alpha}L_{f}h(1-\gamma(1-\sigma))\left(\frac{\Theta}{\theta}\frac{1}{\underline{M}}\right)^{2-\gamma(1-\sigma)}$
$\phantom{aaaaaa}\cdot\frac{\Theta}{\theta^{2}}\left(M_{2}(\mu_{1})^{\frac{1}{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}\right)+\Theta\right)\mathcal{W}_{2}(\mu_{1},\mu_{2}).$
@@
so that
$G_{0}\left(h,p_{0}(x,h,\mu)\right)=A(x)^{1-\sigma}F(x,\mu)^{\gamma(1-\sigma)-1}$
and
$A(x_{1})^{1-\sigma}F(x_{1},\mu)^{\gamma(1-\sigma)-1}<G_{0}(h,p)\leq A(x_{2})^{1-\sigma}F(x_{2},\mu)^{\gamma(1-\sigma)-1}.$
Then
$\big|D_{p}H_{1}(x_{1},h,$ $\mu,p)-D_{p}H_{1}(x_{2},h,\mu,p)\big|$
$=f(h)F(x_{1},\mu)p^{-\frac{1}{1-\eta}}\left(p^{\frac{1}{1-\eta}}-p_{0}(x_{1},h,\mu)^{\frac{1}{1-\eta}}\right)$
$\leq f(h)\frac{\Theta}{\theta}M(\mu)p_{0}(x_{1},h,\mu)^{-\frac{1}{1-\eta}}(1-\gamma)^{\frac{1}{1-\eta}}f(h)^{-1}$
$\phantom{aaaa}\cdot\left(G_{0}(h,p)^{\frac{1}{1-\eta}}-\left(A(x_{1})^{1-\sigma}F(x_{1},\mu)^{\gamma(1-\sigma)-1}\right)^{\frac{1}{1-\eta}}\right)$
$=f(h)F(x_{1},\mu)\left(A(x_{1})^{1-\sigma}F(x_{1},\mu)^{\gamma(1-\sigma)-1}\right)^{-\frac{1}{1-\eta}}$
$\phantom{aaaa}\cdot\left(G_{0}(h,p)^{\frac{1}{1-\eta}}-G_{0}\left(h,p_{0}(x_{1},h,\mu)\right)^{\frac{1}{1-\eta}}\right)$
@@
$\leq L_{f}h\left(\frac{\Theta}{\theta}M(\mu)\right)^{1+\alpha}\underline{A}^{-\beta}$
$\phantom{aaaa}\cdot\left(F(x_{1},\mu)^{-\alpha}\left|A(x_{1})^{\beta}-A(x_{2})^{\beta}\right|+A(x_{2})^{\beta}\left|F(x_{1},\mu)^{-\alpha}-F(x_{2},\mu)^{-\alpha}\right|\right)$
@@
$\begin{cases}\partial_{t}\mu(t)=\frac{1}{2}\epsilon^{2}D^{2}_{xx}\mu(t)+\frac{1}{2}\chi^{2}D^{2}_{hh}\left(h^{2}\mu(t)\right)-D_{x}\left(D_{p}H_{0}(D_{x}V(t,x,h))\mu(t)\right)\\ \phantom{aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa}-D_{h}\left(D_{p}H_{1}(x,h,\mu(t),D_{h}V(t,x,h))\mu(t)\right),\\ \mu(0)=\mu_{0}\end{cases}$ (34)
@@
There exists $T>0$ such that we have existence of a strong solution ($)(\tilde{x}(\cdot),\tilde{h}$($\cdot$)) to (35) on $[0,T]$, which is morevoer pathwise unique.
@@
$\left|\begin{pmatrix}D_{p}H_{0}\left(D_{x}V(t,x_{1},h_{1})\right)-D_{p}H_{0}\left(D_{x}V(t,x_{2},h_{2})\right)\\ D_{p}H_{1}\left(x_{1},h_{1},\mu(t),D_{h}V(t,x_{1},h_{1})\right)-D_{p}H_{1}\left(x_{2},h_{2},\mu(t),D_{h}V(t,x_{2},h_{2})\right)\end{pmatrix}\right|\\ \leq L_{H,N}\left|(x_{1}-x_{2},h_{1}-h_{2})\right|.$
@@
If $\overline{M}(\mu(\cdot))=0$ (so that all the $h$-marginals of $\mu(t)$ equal $\delta_{0}$) then equation (36) clearly has a unique solution. In particular $h(\cdot)$ is then a geometric Brownian motion starting at $h_{0}$, so that $h(t)=0$ $\mathbb{P}$-a.s. for every $t$ if $h_{0}=0$ $\mathbb{P}$-a.s.. Moreover if $h_{0}=0$ $\mathbb{P}$-a.s. then the $\mathbb{P}$-a.s. constant process $h(\cdot)=0$ is a solution to the second equation in (36) for any $\mu(\cdot)$.
Suppose now that $h_{0}>0$ $\mathbb{P}$-a.s. and define
@@
Conversely, take any solution ($)(\hat{x}(\cdot),\hat{h}$($\cdot$)) to (36) with $h_{0}>0$ $\mathbb{P}$-a.s.; the map $t\mapsto(\hat{x}(t),\hat{h}(t))$ is almost surely continuous, so that setting $\tau=\inf\left\{t\geq 0\colon h(t)\leq 0\right\}$ we have $\tau>0$ almost surely. Then for $t>0$ we have, by Ito formula and (31),
@@
Pathwise uniqueness of solutions to (36) when $h_{0}=0$ $\mathbb{P}$-a.s. is a simple consequence of local Lipschitzianity of $h\mapsto D_{p}H_{1}(x,h,\mu,p)$ and boundedness of $D_{h}V$ on $0,T_{1}]\times\mathbb{R}\times\mathbb{R}_{+}$ (see the proof of Lemma 3.6).
If $0<\mu_{0}\left(\mathbb{R}\times\{0\}\right)<1$ (i.e. if $0<\mathbb{P}(h_{0}=0)<1$), we can consider the two conditional laws of ($)(\tilde{x}(\cdot),\tilde{h}$($\cdot$)) with respect to the sets $\{h_{0}=0\}$ and $\{h_{0}>0\}$. By the arguments above, the measure defined by
@@
for every measurable $A\subset C([0,+\infty))$ gives a weak solution to (36); moreover the process ($)(\hat{x}(\cdot),\hat{h}$($\cdot$)) with law $\mathscr{L}$ is pathwise unique, therefore the solution is strong by the Yamada-Watanabe theorem.
This shows that, for every $\mu(\cdot)\in C^{\frac{1}{2}}\left([0,T_{1}];\mathcal{P}_{2}\right)$ such that $\mu(0)=\mu_{0}$ and every $\mathbb{R}\times\mathbb{R}_{+}$-valued random variable $(x_{0},h_{0})$ with law $\mu_{0}$, equation (36) has a unique strong solution on $0,T_{1}]$.
@@
[…]
@@
[…]
$\mathbb{E}$ $\left[\sup_{t\in[0,T]}\hat{h}(t)^{2}\right]$
$\leq 3\mathbb{E}[h_{0}^{2}]+3Tg_{1}^{2}(\overline{M}(\mu(\cdot)))\int_{0}^{T}\mathbb{E}\left[\sup_{s\in[0,t]}h(s)^{2}\right]\mathrm{d}t+3\chi^{2}\mathbb{E}\left[\sup_{t\in[0,T]}\left|\int_{0}^{t}h(s)\mathrm{d}W(s)\right|^{2}\right]$
@@
[…]
@@
If instead $0<\mu_{0}(\mathbb{R}\times\{0\})<1$ then $M(\mu_{0})>0$ and we can again condition on the sets $\{h_{0}=0\}$ and $\{h_{0}>0\}$. As the solution ($)(\tilde{x}(\cdot),\tilde{h}$($\cdot$)) is uniquely determined on each set, well-posedness of (35) follows. ∎
@@
$\int_{\mathbb{R}\times\mathbb{R}_{+}}u(t,x,h)\mu(t,\mathrm{d}x,\mathrm{d}h)-\int_{\mathbb{R}\times\mathbb{R}_{+}}u(t_{0},x,h)\mu(t_{0},\mathrm{d}x,\mathrm{d}h)\\ =\int_{t_{0}}^{t}\int_{\mathbb{R}\times\mathbb{R}_{+}}\frac{\partial u}{\partial t}(r,x,h)\mu(r,\mathrm{d}x,\mathrm{d}h)\mathrm{d}r\\ +\frac{1}{2}\int_{t_{0}}^{t}\int_{\mathbb{R}\times\mathbb{R}_{+}}\frac{1}{2}\mathrm{Tr}\left[g(x,h)g^{\ast}(x,h)D^{2}u(r,x,h)\right]\mu(r,\mathrm{d}x,\mathrm{d}h)\mathrm{d}r\\ +\int_{t_{0}}^{t}\int_{\mathbb{R}\times\mathbb{R}_{+}}e^{\mu}(r,x,h)\cdot Du(r,x,h)\mu(r,\mathrm{d}x,\mathrm{d}h)$
@@
$\int_{\mathbb{R}\times\mathbb{R}_{+}}\phi(x,h)\mu(t,\mathrm{d}x,\mathrm{d}h)=\int_{\mathbb{R}\times\mathbb{R}_{+}}u(0,x,h)\mu_{0}(\mathrm{d}x,\mathrm{d}h);$